1. A debt of $6479.22 is due June 1,2021.what is the value of the obligation on march 1,2016, if money is worth 2% compounded semi-annually?

the value of the obligation is $ ?

(round to the nearest cent as needed . Round all intermediate values to six decimal places as needed)


1
Expert's answer
2022-03-28T16:13:34-0400

The present value is given as

"PV=FV \left( 1 + \frac{ r}{ m } \right) ^ {( - \left( n \times m \right) )}"

Where FV is $6479.22

R is 2%(0.02)

N is 4.25

M is 2.


"PV=\$6479.22 \left( 1 + \frac{ 0.02 }{ 2 } \right) ^ {( - \left( 4.25 \times 2 \right) )}"

"PV=\$ 5953.76"

The money is worth $5953.76